OpenStudy (anonymous):

find all solutions b/w [0, 2pi): cot theta+sq root of 3=0. write in radians in terms of pi

OpenStudy (jdoe0001):

can you factor \( \large cot(\theta)+\sqrt{3}=0 \ ?\)

OpenStudy (anonymous):

i have no clue but its just asking for the solutions so i thought i could simplify to cot theta= - sq 3

OpenStudy (jdoe0001):

that IS the answer :), now just check your Unit Circle for an angle where COT is \(\large -\sqrt{3}\), keeping in mind that COT is negative, thus that means II and IV quadrants only

OpenStudy (anonymous):

:) im not sure which angles those are, thats the issue. could u please tell me?

OpenStudy (jdoe0001):

ok, in your Unit Circle in the II and IV quadrants, look at say, the angles that have a \(\sqrt{3}\) in their SIN and COS values

OpenStudy (jdoe0001):

there are 2 in each quadrant

OpenStudy (jdoe0001):

now, once you find them, keeping in mind that \( \large cot(\theta)=\cfrac{cos(\theta)}{sin(\theta)} \), divide them :) to get a value

OpenStudy (anonymous):

ok i tried that but it still doesnt make sense to me because i got strange decimals.. can u please tell me the answer? i really need help